4. For the collection of capacitors shown below... a. b. find the equivalent capacitance of the combination. find the charge that collects on the 4 μF capacitor.
Answer #1

УНЕ ut IAF 12V 12Y 12V

(a)

4\mu F \: and \: 8\mu F are in series, so

C'=\frac {4\times 8}{4+8}\: \mu F= 2.67\:\mu F

C' \: and\: 1\mu F are in parallel, so

C_{eq}= (2.67+1)\:\mu F

C_{eq}= 3.67\:\mu F

(b)

V_{1\mu F}=V_{C'}= 12\: V

Q_{4\mu F}= C'V_{C'}=32.04\: \mu C